OpenStudy (anonymous):

2x + 5y = −3 −15x − 5y = −10 Solve the following system of inequalities using elimination, substitution or graphing method.

OpenStudy (anonymous):

just add the two, you will consquently get rid of the y variable, then just solve for x

OpenStudy (anonymous):

I do not understand what you mean my friend :/ Is it: -15x+2x=-10? makes no sense since that is addition, you are not eliminating nor substituting.

OpenStudy (anonymous):

if you add the two equations you are elimintang the y variable

OpenStudy (anonymous):

2x + 5y = −3 −15x − 5y = −10 ---------------- -13x=-13

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Then sub x=1 into either of the orginal inequalities and solve for y, my friend

OpenStudy (anonymous):

2x + 5y = −3 + −15x − 5y = −10 _______________ -13x+0=-7 -13x=-7 -7/13=x 2(-7/13)+5y=-3 -14/13+5y=-3 5y=-25/13 y=-5/13 thats what I have :)

OpenStudy (anonymous):

2x + 5y = −3 + −15x − 5y = −10 _______________ -13x+0=-13 -13x=-13 1=x 2+5y=-3 5y=-5 y=-1 correct answer

OpenStudy (anonymous):

isnt that what i had x=1

OpenStudy (anonymous):

yes the first one i did was wrong. i forgot about the signs :P again